OpenStudy (anonymous):

if csc( weird squiggly thing) = 13/5 then what would sec sin cos and tan = ?

OpenStudy (amistre64):

design a rt triangle with sides corresponding to the definition of csc, and then you can determine the missing side to fill in the other trig ratios

OpenStudy (amistre64):

or you can do a pythag method that doesnt rely on geometry

OpenStudy (anonymous):

so then 13 is the long side 5 is the short side and when added together its 18?

OpenStudy (amistre64):

almost, its good to remember that sin and csc are flips of each other so if csc = 13/5, then sin = 5/13 sin = over/hyp